An RLC series circuit has a 2 kΩ resistor, a 373 mH inductor, and a 29.7 LaTeX: \muμF capacitor. If this is connected to a 391 Volt power supply, what will the rms current be at 344 Hz? Express your answer in mA.
Answer #1

formula for Rms corrent, Irms = Voms 2= √ R²+ (XL-Xc)? XL = 27142= 2T (344)(373x10-3)= 80622. * 2T (344)(a97X106) – 15.54 r.